OpenStudy (kawaiichan7678):

Hi I'm new to OpenStudy. Can someone help me with these questions? thanks! <3

OpenStudy (kawaiichan7678):

2.|y+4|<1 A.-5<y<-3 B.-3<y<5 C.-4<y<1 D.1<y<4

likeabossssssss (likeabossssssss):

welcome to os im neew to lol :)

OpenStudy (kawaiichan7678):

thanks lol I wasn't expecting someone to respond so fast

OpenStudy (kawaiichan7678):

Can someone help me? I really don't understand 3 of these questions

OpenStudy (shansha1):

i think its b

OpenStudy (kawaiichan7678):

okay thank you and this one? 3.|2t|-5=7 A.t=1 or -1 B. t=6 or -6 C. t= 10 or -10 D.t=12 or-12

OpenStudy (kittiwitti1):

Direct answers are not allowed. ☻ \[|2t|-5=7\]\[|2t|\cancel{-5}\cancel{+5}=7-5\]\[|2t|=7-5=?\]You may simplify the value after that. Remember, absolute values have both a negative and positive value!\[\large{|ab|=\pm ab}\]

OpenStudy (kawaiichan7678):

Thank you I don't quite understand yet but thank you so much! <3

OpenStudy (kittiwitti1):

What don't you understand?

OpenStudy (shanedaigle):

I think its B

OpenStudy (kawaiichan7678):

I've never done any of this before I went into this school late and they put me directly into this lesson

OpenStudy (shanedaigle):

Hay I'm not sure if its the answer or not its just my best guess

OpenStudy (kawaiichan7678):

thanks though it's better than me guessing

OpenStudy (shanedaigle):

I still think its B

OpenStudy (kawaiichan7678):

Okay thank you

OpenStudy (kawaiichan7678):

|a|-3/4=-5/8 A.1/8 or -1/8 B.7/8 or -7/8 C.1 3/8 or -1 3/8 D. no solution

OpenStudy (kawaiichan7678):

that's the last one <3

OpenStudy (kawaiichan7678):

it's A right?

OpenStudy (kawaiichan7678):

nvm I got it! 1.B 2.A 3.B 4.A Thanks so much guys! <3

OpenStudy (sachintha):

You go the 3rd one wrong. Try to work it out from kittiwitti's answer. :)

OpenStudy (shanedaigle):

??????

OpenStudy (shanedaigle):

I did # 3 out and got B as the answer.

OpenStudy (sachintha):

Shane you are correct. Pardon me. I just pondered over kitti's reply and didn't notice that he did a mistake.

OpenStudy (shanedaigle):

:)

OpenStudy (kittiwitti1):

Oh, whoops. I seem to have made a signage error. My bad